Demonstration of ALF

Catarina Coquand  Lena Magnusson

Chalmers Technical University Inst. for Computer Science Göteborg, Sweden

            E-mail: {catarina,lena}@cs.chalmers.se

We refer to our abstracts [Coq] and [Mag] in these proceedings for further information about the demonstrated system.

References

[Coq] Catarina Coquand. Introduction to ALF — a background. Chalmers Technical University, Inst. for Computer Science, Göteborg, Sweden. E-

      mail: catarina@cs.chalmers.se.

[Mag] Lena Magnusson. Introduction to ALF — an interactive proof editor. Chalmers Technical University, Inst. for Computer Science, Göteborg,

      Sweden. E-mail: lena@cs.chalmers.se.